find the solution to each system by the substitution method
2x-5y=-1
x-3y=-3
next problem
10x+3y=8
2x+y=2
email to send the answers cgividen0817@yahoo.com Thank you

Answer by mananth(16946)

2 x -5 y = -1 .............1
-3 y = -3 -1 x ---------2
/ -3 =
y = 1 + 0.33 x
Substitute y in (1)
2 x -5 ( 1 + 0.33 x ) = -1
2 x -5 + -1.67 x = -1
2 x -1.67 x = -1 5
0.33 x = 4
/ 0.33
x= 12
Plug the value of x in (1)
2 x -5 y = -1
2 * 12 -5 y = -1
24 -5 y = -1
-5 y = -25
/ -5
y= 5
-----------------
10 x + 3 y = 8 .............1
1 y = 2 -2 x ---------2
/ 1 =
y = 2 -2 x
Substitute y in (1)
10 x + 3 ( 2 -2 x ) = 8
10 x + 6 + -6 x = 8
10 x -6 x = 8 -6
4 x = 2
/ 4
x= 0.5
Plug the value of x in (1)
10 x + 3 y = 8
10 * 0.5 + 3 y = 8
5 + 3 y = 8
3 y = 3
/ 3
y= 1
m.ananth@hotmail.ca